Hi
Bob had some money left after spending $9 on 4 files and 6 pens. If he were to buy one more file, he would be short 30cents. If he were to buy one more pen he would have 70 cents left.
What was the cost of each file.
How much money did he have at first.
Thanks

Answer by ikleyn(52776)

.

            This problem is tricky:  I made several attempts before I came to right setup.


Let F be the price for each file;  P be the price for each pen, and

let X be the amount of money that Bob had initially.


Then, based on the problem's description, we can write these THREE equations for three unknowns


    4F + 6P = 900     cents     (1)     (he spent $9 on 4F and 6P)

    5F + 6P = X + 30  cents     (2)     (he would be short 30 cents)

    4F + 7P = X - 70  cents     (3)     (he would be 70 cents left)



        The setup is just completed. 
        As I said at the beginning, it is tricky.
        But the solution of equations is simple.



To solve them, first subtract equation (3) from equation ((2).  You will get

    F - P   =     100  cents.   (4)


Next, express  F = P + 100 from (4), and substitute it into equation (1).  You will get

    4(P + 100) + 6P = 900

        10P         = 900 - 400 = 500

          P                     = 500/10 = 50.


Next, from equation (1),  4F + 6*50 = 900,  4F = 900 - 300 = 600,  F = 600/4 = 150.


Thus, the price for one file is $1.50;  the price for one pen is  P = F - 100 = 150-100 = 50 = $0.5  (from equation (4))

and  X = 5F + 6P - 30 = 5*150 + 6*50 - 30 = 1020 cents = $10.20.  (from equation (2) )


ANSWER.  The price for each file is  $1.50.  Bob had  $10.20  originally.

Solved.     //   All questions are answered.